OpenStudy (anonymous):

10x-y+2z=-17, x+2y-z=4, 2x+2y-3z=9

OpenStudy (tkhunny):

10x-y+2z=-17 x+2y-z=4 2x+2y-3z=9 What have you tried? If it were me, I'm multiply the first by 2 20x-2y+4z=-34 Then add it to the other two equations. What then?